OpenStudy (anonymous):

Solve x to the third power = 1 over 8. 1 over 2 ±1 over 2 1 over 4 ±1 over 4

OpenStudy (anonymous):

Any Help would be generous

OpenStudy (anonymous):

\[x^3=\frac{1}{8}\] means \(x=\frac{1}{2}\) since \[\left(\frac{1}{2}\right)^3=\frac{1}{8}\]

OpenStudy (anonymous):

The question make signs are these ±

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw